# Unveiling the Hidden Structure of Self-Attention via Kernel Principal Component Analysis

### v1

@@ A.2 Robustness under Adversarial Attacks
We evaluate the robustness of our method against adversarial attacks using CleverHans [44] and AutoAttack [14]. All attacks are executed on ImageNet-1K’s validation set, and each model is evaluated on the whole set. In particular, we use untargeted, white box attacks such as PGD, FGSM, SLD and CW. In addition, we provide results on gradient-free black box attack, SPSA, diverse AutoAttack, in the standard setting and a simple noise-adding attack. AutoAttack consists of untargeted APGD-CE, targeted APGD-DLR, targeted Fast Adaptive Boundary (FAB) and Square Attack. For further justification of the benefit of our method, we evaluate a variant of our model, RPC-SymViT (6iter/layer1) solely on the black-box Square attack and show that we are robust against black box attacks as well. This is provided in Appendix E.4. We use a perturbation budget of $\epsilon=1/255$ with the $l_{\infty}$ norm to manipulate the images and evaluate each model with an incremental increase of $1/255$ in perturbation for PGD and FGSM. In SPSA, we run 40 steps per attack with a perturbation budget of $\epsilon=0.1$. PGD attack uses a step size of $\alpha=0.15$ for 20 steps, where at each step the image is adjusted slightly to maximize the model’s loss. Similarly, FGSM does this for a single step. For each attack, we report the top-1 and top-5 accuracy of the model on the distorted dataset with the maximum perturbation budget in Table 2 and across all different perturbations in Figure 3.
@@ Appendix D Principal Component Pursuit
and solves a sequence of optimization problems. We iterate through ${\bm{S}}_{k+1}=\arg\min_{\bm{S}}l({\bm{L}}_{k},{\bm{S}},{\bm{Y}}_{k})$ and ${\bm{L}}_{k+1}=\arg\min_{\bm{L}}l({\bm{L}},{\bm{S}}_{k+1},{\bm{Y}}_{k})$ before updating the Lagrange multiplier matrix ${\bm{Y}}_{k+1}={\bm{Y}}_{k}+\mu({\bm{M}}-{\bm{L}}_{k+1}-{\bm{S}}_{k+1})$. The advantage of the algorithm is that it turns a complicated optimisation problem into sub problems that have straightforward and efficient solutions. Without much difficulty we can show that:
@@ E.1 Reparameterization of Self-Attention
Let ${\bm{A}}_{sym}=[a_{ij}]$ be the symmetric softmax attention matrix. Then, from Remark 3, ${\bm{S}}$ has the following form and we multiply the numerator and denominator by $1/K({\bm{k}}_{j},{\bm{k}}_{j^{\prime}})$ to obtain a much more convenient expression.
@@ E.5 Results on ADE20K Image Segmentation
E.6 Results on RPC-SymViT-small
To show that RPC-Attention is not limited to small-scale models, we experiment on a larger model, SymViT-small with 12 transformer layers, 6 heads per layer and a hidden dimension of 384. We train a SymViT-small baseline on ImageNet-1K for 300 epochs and finetune that model for 100 epochs with RPC-Attention in the first layer using 4 iterations with the hyperparameters $\lambda=4$ and $\mu=ND/4\|{\bm{K}}\|_{1}$. We refer to this model as RPC-SymViT-small (4iter/layer1). We compare our RPC-SymViT-small with the baseline on the same standard robustness benchmarks as before, ImageNet-C, ImageNet-A, ImageNet-O and ImageNet-R as well as on white box attacks PGD and FGSM with the highest perturbation budgets. These results are in Table 7 and 8 respectively and show that RPC-Attention is also effective in SymViT-small.
Table 7: Top-1, Top-5 accuracy (%) , mean corruption error (mCE), and area under the precision-recall curve (AUPR) of RPC-SymViT-small and SymViT-small on clean ImageNet-1K data and popular standard robustness benchmarks for image classification. RPC-SymViT-small ($n$iter/layer1) applies $n$ PAP iterations only at the first layer.
Model IN-1K IN-R IN-A IN-C IN-O
SymViT-small (baseline) 78.72 94.39 38.33 17.43 54.32 58.24 20.06
RPC-SymViT-small (4iter/layer1) 78.84 94.49 38.08 17.52 54.48 58.00 20.30
Table 8: Top-1 and Top-5 accuracy (%) of RPC-SymViT-small and SymViT-small on PGD and FGSM attacked ImageNet-1K validation data with the highest perturbation budget. RPC-SymViT ($n$iter/layer1) applies $n$ PAP iterations only at the first layer.
@@ E.6 Results on RPC-SymViT-small
SymViT-small (baseline) 41.47 71.53 8.35 19.99
RPC-SymViT-small (4iter/layer1) 41.94 71.59 8.94 20.70
@@ E.7 Results on RPC-Attention in FAN
E.8 Computational Efficiency
@@
E.9 Results on Robust Asymmetric Attention
@@
In this section, we report the results of extending the RPC-SymViT model to the asymmetric attention. However, as the PAP Algorithm. 1 is not designed for multiple data matrices, it is not as effective in the asymmetric case. We implement two variations of the algorithm in an asymmetric attention ViT-tiny with 12 layers. In the 4iter/layer1 version, similar to the symmetric attention case, we run 4 iterations of the algorithm only in the first layer of the model by replacing $\text{Softmax}({\bm{K}}-{\bm{S}}_{k+1}-\mu^{-1}{\bm{Y}}_{k},{\bm{K}}-{\bm{S}}_{k+1}-\mu^{-1}{\bm{Y}}_{k})$ with $\text{Softmax}({\bm{Q}},{\bm{K}}-{\bm{S}}_{k+1}-\mu^{-1}{\bm{Y}}_{k})$. For the second version, labeled Sym/Asym, we run the algorithm for 4 iterations in a symmetric attention mechanism in the first layer, followed by the standard asymmetric attention in layers 2 to 12. We compare these RPC-AsymViT models to the asymmetric softmax attention ViT.

### v2

@@ A.2 Robustness under Adversarial Attacks
We evaluate the robustness of our method against adversarial attacks using CleverHans [59] and AutoAttack [18]. All attacks are executed on ImageNet-1K’s validation set, and each model is evaluated on the whole set. In particular, we use untargeted, white box attacks such as PGD, FGSM, SLD and CW. In addition, we provide results on gradient-free black box attack, SPSA, diverse AutoAttack, in the standard setting and a simple noise-adding attack. AutoAttack consists of untargeted APGD-CE, targeted APGD-DLR, targeted Fast Adaptive Boundary (FAB) and Square Attack. For further justification of the benefit of our method, we evaluate a variant of our model, RPC-SymViT (6iter/layer1) solely on the black-box Square Attack and show that we are robust against black box attacks as well. This is provided in Appendix E.4. We use a perturbation budget of $\epsilon=1/255$ with the $l_{\infty}$ norm to manipulate the images and evaluate each model with an incremental increase in perturbation for PGD and FGSM. In SPSA, we run 40 steps per attack with a perturbation budget of $\epsilon=0.1$. PGD attack uses a step size of $\alpha=0.15$ for 20 steps, where at each step the image is adjusted slightly to maximize the model’s loss. Similarly, FGSM does this for a single step. For each attack, we report the top-1 and top-5 accuracy of the model on the distorted dataset with the maximum perturbation budget in Table 2 and across all different perturbations in Figure 3.
@@ Appendix D Principal Component Pursuit
and solves a sequence of optimization problems. We iterate through setting ${\bm{S}}_{k+1}=\arg\min_{\bm{S}}l({\bm{L}}_{k},{\bm{S}},{\bm{Y}}_{k})$ and ${\bm{L}}_{k+1}=\arg\min_{\bm{L}}l({\bm{L}},{\bm{S}}_{k+1},{\bm{Y}}_{k})$ before updating the Langrange multiplier matrix ${\bm{Y}}_{k+1}={\bm{Y}}_{k}+\mu({\bm{M}}-{\bm{L}}_{k+1}-{\bm{S}}_{k+1})$. The advantage of the algorithm is that it turns a complicated optimisation problem into sub problems that have straightforward and efficient solutions. Without much difficulty we can show that:
@@ E.1 Reparameterization of Self-Attention
Let ${\bm{A}}_{sym}=[a_{ij}]$ be the softmax attention matrix. Then, from Remark 3, ${\bm{S}}$ has the following form and we multiply the numerator and denominator by $1/K({\bm{k}}_{j},{\bm{k}}_{j^{\prime}})$ to obtain a much more convenient expression.
@@ E.5 Results on ADE20K Image Segmentation
E.6 Results on RPC-SymViT-base
To show that RPC-Attention is not limited to small-scale models, we conduct experiments on a larger model, SymViT-base with 12 transformer layers, 12 heads per layer, and a hidden dimension of 768. We train a SymViT-base with RPC-Attention in all layers using 2 iterations on ImageNet-1K for 300 epochs. We refer to this model as RPC-SymViT-base (2iter/all-layer). We compare our RPC-SymViT-base with the baseline on the same standard robustness benchmarks as before, ImageNet-C, ImageNet-A, and ImageNet-O, as well as on white box attacks PGD and FGSM with the highest perturbation budgets. These results are in Table 7 and 8 respectively and show that RPC-Attention is also effective in SymViT-base.
Table 7: Top-1, top-5 accuracy (%) , mean corruption error (mCE), and area under the precision-recall curve (AUPR) of RPC-SymViT-base and SymViT-base on clean ImageNet-1K data and popular standard robustness benchmarks for image classification. RPC-SymViT-base ($n$iter/all-layer) applies $n$ PAP iterations at all layers.
Model IN-1K IN-A IN-C IN-O
Top-1 $\uparrow$ Top-5 $\uparrow$ Top-1 $\uparrow$ Top-1 $\uparrow$ mCE $\downarrow$ AUPR $\uparrow$
SymViT-base (baseline) 80.62 94.78 24.03 58.88 52.57 23.96
RPC-SymViT-base (2iter/all-layer) 80.72 94.82 24.97 59.29 52.00 26.01
Table 8: Top-1 and top-5 accuracy (%) of RPC-SymViT-base and SymViT-base on PGD and FGSM attacked ImageNet-1K validation data with the highest perturbation budget. RPC-SymViT-base ($n$iter/all-layer) applies $n$ PAP iterations at all layers.
@@ E.6 Results on RPC-SymViT-base
SymViT-base (baseline) 12.11 25.55 53.61 78.05
RPC-SymViT-base (2iter/all-layer) 13.41 26.68 54.68 78.06
@@ E.7 Results on RPC-Attention in FAN
E.8 Results on downstream Natural Language Understanding tasks
Table 11: Train and validation accuracy (%) of RPC-SymSMoE, SymSMoE, RPC-SMoE (asymmetric) and SMoE (asymmetric) on SST-2 and IMDB downstream tasks when pre-trained on WikiText-103. RPC is only implemented during fine-tuning, not pre-training and runs for 2 iterations for all layers in each model.
Model SST-2 IMDB
Train Acc $\uparrow$ Valid Acc $\uparrow$ Train Acc $\uparrow$ Valid Acc $\uparrow$
SymSMoE (baseline) 63.12 69.20 - -
RPC-SymSMoE 74.48 76.27 - -
SMoE (baseline) 53.20 69.38 64.12 65.74
RPC-SMoE 58.87 70.63 74.36 73.96
Table 12: Validation and test accuracy of RPC-Attention implemented in a pre-trained transformer language model (RPC-LM) during fine-tuning versus the baseline transformer language model (Baseline-LM). We fine-tune both models on the 5-class Stanford Sentiment Treebank (SST-5) task with 2 iterations for all layers in the RPC model.
Model Valid Acc $\uparrow$ Test Acc $\uparrow$
Baseline-LM 46.51 49.23
RPC-LM 48.68 50.36
We conduct additional experiments on several downstream Natural Language Understanding tasks to illustrate the effectiveness of RPC-Attention during fine-tuning as well. The Stanford Sentiment Treebank v2 (SST-2) and IMDB Sentiment Analysis (IMDB) tasks are binary classifications where the goal is to determine if sentences have positive or negative sentiments. We use 2 baseline Sparse Mixture of Experts (SMoE) models, one with symmetric attention (SymSMoE) and one with asymmetric attention (SMoE), pre-trained on WikiText-103 without RPC, then fine-tuned with (RPC-SymSMoE/RPC-SMoE) and without RPC-Attention for comparison. Their results can be found in Table 11. We observe that RPC-models outperform the baseline models significantly on these tasks.
On the 5-class sentiment classification task, Stanford Sentiment Treebank (SST-5), we use a pre-trained transformer language model from the NAACL 2019 tutorial on “Transfer Learning in Natural Language Processing” for fine-tuning. Their publicly available code can be found at https://github.com/huggingface/naacl_transfer_learning_tutorial. The objective of the task is to determine if the sentences have negative, somewhat negative, neutral, somewhat positive, or positive sentiments. We implement RPC-Attention during fine-tuning only (RPC-LM) and compare the results with the baseline model (Baseline-LM) on SST-5 in Table 12. As can be seen from the table, our RPC-Attention is applicable to pre-trained language models and performs significantly better than the baseline. Hence, RPC-Attention is highly effective in downstream natural language understanding tasks and versatile in its application.
E.9 Computational Efficiency
@@
E.10 Results on Robust Asymmetric Attention
@@
In this section, we report the results of extending the RPC-SymViT model to the asymmetric attention. However, as the PAP Algorithm. 1 is not designed for multiple data matrices, it is not as effective in the asymmetric case. We implement two variations of the algorithm in an asymmetric attention ViT-tiny with 12 layers. In the 4iter/layer1 version, similar to the symmetric attention case, we run 4 iterations of the algorithm only in the 1st layer of the model by replacing $\text{Softmax}({\bm{K}}-{\bm{S}}_{k+1}-\mu^{-1}{\bm{Y}}_{k},{\bm{K}}-{\bm{S}}_{k+1}-\mu^{-1}{\bm{Y}}_{k})$ with $\text{Softmax}({\bm{Q}},{\bm{K}}-{\bm{S}}_{k+1}-\mu^{-1}{\bm{Y}}_{k})$. For the second version, labeled Sym/Asym, we run the algorithm for 4 iterations in a symmetric attention mechanism in the 1st layer, followed by the standard asymmetric attention in layers 2 to 12. We compare these RPC-AsymViT models to the asymmetric softmax attention ViT.
@@ Appendix F Further Discussion on Related Works
Hamburger in [25] models the global context discovery as the low-rank recovery of the input tensor and solves it via matrix decomposition. Both Hamburger and our Attention with Robust Principal Components (RPC-Attention) try to recover clean signal subspaces via computing a low-rank approximation of a given matrix. The key differences between our RPC-Attention and Hamburger are: (1) Our RPC-Attention finds a low-rank approximation of the key matrix while Hamburger finds a low-rank approximation of the input matrix, and (2) Our RPC-Attention models the corruption by a sparse matrix while Hamburger does not enforce this condition. The entries of this sparse corruption can have an arbitrarily large magnitude and help model grossly corrupted observations in which only a portion of the observation vector is contaminated by gross error. Numerous critical applications exist where the data being examined can be naturally represented as a combination of a low-rank matrix and a sparse contribution, such as video surveillance, face recognition, and collaborative filtering [11].
[89] derives each component in a Graph Neural Network (GNN) from the unfolded iterations of robust descent algorithms applied to minimizing a principled graph regularized energy function. In particular, propagation layers and nonlinear activations implement proximal gradient updates, and graph attention results from iterative reweighted least squares (IRLS). While this is an interesting approach, it has not been extended to explaining the architecture of transformers, including self-attention, yet. In contrast, our kernel principal component analysis (kernel PCA) allows us to derive self-attention in transformers, showing that the attention outputs are projections of the query vectors onto the principal components axes of the key matrix in a feature space.
[2] and [6] implement each layer as an optimization and fixed-point solver, respectively. In particular, an OptNet layer in [2] solves a quadratic program, and a Deep Equilibrium layer in [6] computes the fixed point of a nonlinear transformation. Different from these layers, our RPC-Attention solves a Principal Component Pursuit - a convex program. Also, both OptNet layer in [2] and Deep Equilibrium layer in [6] do not shed light on the derivation and formulation of self-attention, which our kernel PCA framework does.